Step by step on how to solve: 8x - 2 = -9 + 7x
Answer by RicaMath(8)

To solve this problem, you need to isolate the x variable.


1) Subtract 7x from both sides.

8x - 2 - 7x = -9 + 7x - 7x
x - 2 = -9


2) Add 2 to each side.

x - 2 + 2 = -9 + 2
x = -7


You can always double check your work by plugging the answer back into the original problem:

8x - 2 = -9 + 7x
8(-7) - 2 = -9 + 7(-7)
-56 - 2 = -9 - 49
-58 = -58

This is a true statement, therefore the answer is indeed x = -7.